OpenStudy (anonymous):

Given the function f described by f(x) =9x, find each of the following: (a) f(-2) (b) f(4) (c) f(26) Answer would be f(-2) = ??

OpenStudy (anonymous):

f(x) means function of a variable x. f(-2) means function of -2. so substitute x = -2 in the expression and find the value of f(x)

OpenStudy (anonymous):

So... f(-2)=9*(-2) f(4)=9*4 f(26)=9*26 Just simplify the right side of those equations and there are your answers :)

OpenStudy (anonymous):

I don't know how to do that. It is just so confusing for me.

OpenStudy (anonymous):

okay f(x) = 9x means f(x) is function of x such that for a given value of x, the value of the function is 9 times that value of x. so f(-2) means the value of the function at x = -2 so that is 9 times -2 = -18.

OpenStudy (anonymous):

so f(-2) = 9?

OpenStudy (anonymous):

No, f(-2)=9*(-2), or -18 Similarly, f(4)=9*4, or 36.

OpenStudy (anonymous):

read what I posted.

OpenStudy (anonymous):

I am sorry I am so stupid. That's why I am "dumbunny". So f(-2) = 36?

OpenStudy (anonymous):

no. read what I posted. f(-2) is 9 times -2 = -18. the function takes a value x and multiplies it by 9. that is all it does. so the function is like a magnifying glass. you put an ant on one side of the glass and it appears 9 times larger at the other end.

OpenStudy (anonymous):

so if you put 4 at one end, it makes it 9 times larger, that is, it makes it appear as 36. so f(4) = 36.

OpenStudy (anonymous):

so f(-2) = 18 or -18? Right?

OpenStudy (anonymous):

no. f(-2) = -18, not +18

OpenStudy (anonymous):

no...18

OpenStudy (anonymous):

so it would be -18 then

OpenStudy (anonymous):

That was right. thank you. So how about if f(4)=??

OpenStudy (anonymous):

yes. do you get what a function is? suppose you have f(x) = x+3 what this function f(x) does is, takes a variable x and adds 3 to it. so if you have x = 2 f(x) = take 2 and add 3 to it. that is, f(x) = 2 +3 = 5 get it? function means just that. perform some function on an input and give an output.

OpenStudy (anonymous):

so f(x) = 9x means take a variable x and multiply it by 9. so f(4) means take 4 and multiply it by 9.

OpenStudy (anonymous):

NO, unfortunately I don't get this at all. I am a psych major and algebra is not my forte...

OpenStudy (anonymous):

f(4)=?

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

36 was right...how about f(26) =?

OpenStudy (anonymous):

I am getting there.

OpenStudy (anonymous):

write f(26) as a sentence in plain english the same way I wrote it for f(4)

OpenStudy (anonymous):

so would it be 4 x 26?

OpenStudy (anonymous):

or 9 x 26?

OpenStudy (anonymous):

you tell me.

OpenStudy (anonymous):

first write it in plain english. what does the function f(x) do?

OpenStudy (anonymous):

Oh don't tease me...

OpenStudy (anonymous):

I think it would be 9 x 26

OpenStudy (anonymous):

okay. why do you think that?

OpenStudy (anonymous):

I don't know. Because it makes more sense?

OpenStudy (anonymous):

why does it make more sense? Did you read what I posted about this function. what does the function f(x) = 9x do? it takes a variable x and _______ ?

OpenStudy (anonymous):

4 x 26 makes more sense now that I think about it.

OpenStudy (anonymous):

why? can you stop guessing and actually read what I posted earlier? It is better to understand what is going on than to guess.

OpenStudy (anonymous):

I do not get functions at all.

OpenStudy (anonymous):

did you read what I posted or not? Simple question. yes or no?

OpenStudy (anonymous):

I am sorry, I just do not get it. That is why I am guessing.

OpenStudy (anonymous):

yes. do you get what a function is? suppose you have f(x) = x+3 what this function f(x) does is, takes a variable x and adds 3 to it. so if you have x = 2 f(x) = take 2 and add 3 to it. that is, f(x) = 2 +3 = 5 get it?

OpenStudy (anonymous):

YES, I did read what you posted. I just don't understand it.

OpenStudy (anonymous):

okay. now we are getting somewhere. What do you not understand about it?

OpenStudy (anonymous):

Pretty much the whole concept. It just seems to go right over my head.

OpenStudy (anonymous):

Give me psych, biology, history, etc., and I can run circles around anyone, but when it comes to algebra I am a complete moron.

OpenStudy (anonymous):

http://www.twiddla.com/579395 join me there. I will explain the concept.

OpenStudy (anonymous):

unless you don't want to learn it. let me know either way.

OpenStudy (anonymous):

hello? are you there? I do not want to be wasting my time if you are not interested in the answer. so let me know.

OpenStudy (anonymous):

Yes I am here..sorry, had to let my dog out